the sum of 2 integers is 36. the difference of the same 2 numbers is 4. find the integers.
Answer by WannabeCAgirl83(35)


1st integer = x
2nd integer = x+4

1st integer + 2nd integer = 36:
x + x+4 = 36

x + x+4 = 36 | -4 on both sides
x + x = 32 | re-arranging
2x = 32 | :2
x = 16

1st integer = 16
2nd integer = 20 (16+4)